OpenStudy (anonymous):

4. Write a direct variation equation that relates x to y. Then solve. Show both the equation and the solution. If y = 8 when x = 4, find y when x = 5.(Points : 2) 5. Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2).(Points : 1) y = 3x - 5 y = 3x - 2 y = 3x + 1 y = 3x + 3

OpenStudy (blurbendy):

8 / 4 = 2 y / 5 = 2 10 / 5 = 2 y = 10 5) y = 3x -5 because if you put 1 into x, you get -2